# Doc2Chart: Intent-Driven Zero-Shot Chart Generation from Documents

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated strong capabilities in transforming text descriptions or tables to data visualizations via instruction-tuning methods. However, it is not straightforward to apply these methods directly for a more real-world use case of visualizing data from long documents based on user-given intents, as opposed to the user pre-selecting the relevant content manually. We introduce the task of intent-based chart generation from documents: given a user-specified intent and document(s), the goal is to generate a chart adhering to the intent and grounded on the document(s) in a zero-shot setting. We propose an unsupervised, two-staged framework in which an LLM first extracts relevant information from the document(s) by decomposing the intent and iteratively validates and refines this data. Next, a heuristic-guided module selects an appropriate chart type before final code generation. To assess the data accuracy of the generated charts, we propose an attribution-based metric that uses a structured textual representation of charts, instead of relying on visual decoding metrics that often fail to capture the chart data effectively. To validate our approach, we curate a dataset comprising of 1,242 $<$intent, document, charts$>$ tuples from two domains, finance and scientific, in contrast to the existing datasets that are largely limited to parallel text descriptions/ tables and their corresponding charts. We compare our approach with baselines using single-shot chart generation using LLMs and query-based retrieval methods; our method outperforms by upto $9$ points and $17$ points in terms of chart data accuracy and chart type respectively over the best baselines.
@@ 1 Introduction
Advancements in LLMs Brown et al. (2020); Touvron et al. (2023) have enabled recent efforts in generating high-quality statistical charts from user-given text descriptions or tables Wang et al. (2023); Han et al. (2023); Maddigan and Susnjak (2023); Zhang et al. (2024); Tian et al. (2024); Zadeh et al. (2024). However, these existing approaches to automatic chart generation typically assume that users will manually provide tables or textual descriptions to be transformed into charts. Consequently, these methods cannot directly generate charts from raw documents based on specific user intents or queries. Further, many previous methods rely on instruction-tuning techniques that necessitate a substantial volume of labeled data (in the order of 10s of 1000s) for effective training. Obtaining such large corpora of charts paired with their corresponding intents is a tedious and time-consuming process, particularly for long documents (spanning $>$ 5 pages), which are the primary focus of our work.
@@
This paper makes four main contributions: (1) We introduce the novel task of generating charts from documents based on specific intents, to mimic real-world use cases for chart generation. (2) We propose a training-free, multi-stage framework that leverages LLMs to first execute an intent-guided iterative data extraction and refinement process from documents, and subsequently to select an appropriate chart type through a heuristic-guided deliberation, before constructing the final chart. (3) We propose a chart evaluation metric to assess the faithfulness of the generated chart data when compared to either a reference chart/ table (reference-based setting) or source document itself (reference-free setting). (4) We curate a dataset consisting of 1,242 $<$intent, document, chart$>$ tuples from financial reports and scientific articles for this task evaluation.
@@ 2.3 Chart Evaluation
Figure 2: Sample annotations for a given SEC document and its constituent data tables. An intent and a corresponding chart are provided, along with the data table supporting the chart and a description summarizing the chart key takeaways. Other metadata such as page number and tables used from the source document are also provided. We use the page numbers to truncate the documents to include ±$\pm 5$ pages while using the source documents with the chart generation methods, to avoid context length issues.
@@ 3.1 Dataset Curation Overview
For the source documents, we consider two domains, namely finance and scientific. We use the U.S. Securities and Exchange Commission (SEC) 10-K filings [^1] https://www.sec.gov/edgar.shtml that are publicly available on the EDGAR website, and academic papers from *ACL conferences from the SciDuet dataset Sun et al. (2021) respectively. We scrape the HTML documents for 1,000 SEC 10-K filings, and consider all the 1,088 papers from SciDuet in the PDF form. We parse these HTML and PDF documents [^2] BeautifulSoup, https://developer.adobe.com/document-services/docs/overview/pdf-extract-api/ to obtain the tables in them separately in spreadsheets. From among the SEC filings, we first filter documents that contain table(s) with size $>$7x7, and then pick the top 100 ones that have the maximum number of tables in them. As most of the numeric content that can be visualized is present in tables in these documents, we use those that densely contain them. Since academic papers do not always contain very large tables, we take the top 120 documents that have the maximum number of tables in them (and relax the size constraints). Since these documents contain multiple tables, and each table can convey several insights, we obtain multiple intents from each of them, to reduce the number of documents that are to be ingested and the overall cognitive load. Table 1 provides the source document details for the two domains. Figure 2 provides an overview of the data curation process.
@@ 5.2 Evaluation Metrics
[…]
@@ A.1 Qualitative examples
{mdframed}
[backgroundcolor=gray!10, roundcorner=5pt] Task:
@@
Input:
  • User Intent: {intent}
  • Content: {content}
  • Optional Feedback (if available): {optional_feedback_section}
  • Output Format Schema: {output_format}
Instructions:
  1. 1.
    Carefully read the User Intent.
  2. 2.
    Internal Thought Process (Mentally follow these steps):
    • Decompose: Break down the intent into specific data points, labels, categories, and title.
    • Locate: Scan the content for exact data matching the above.
    • Extract & Structure: Collect and format data strictly according to the schema.
  3. 3.
    Extract relevant data points: $(x,y,\text{category})$, axis labels, and chart title.
  4. 4.
    If feedback is provided: Focus on fixing issues like missing elements or ignored sections. Adjust your decomposition and extraction accordingly.
  5. 5.
    Output must follow the JSON schema exactly. Keep numeric formats consistent.
  6. 6.
    Output only the JSON object. Do not include explanations or markdown like ‘‘‘json.
Example Output Format:
@@
{mdframed}
[backgroundcolor=gray!10, roundcorner=5pt] Task: Validate the extracted chart data against the source content and user intent. Determine if re-extraction is necessary or if only minor refinements are needed.
Input:
  • Original Intent: {intent}
  • Source Content: {content}
  • Extracted Chart Data: {extracted_data} // JSON object from the extraction step
  • Expected Schema: {output_format}
Validation Checks to Perform:
  1. 1.
    Intent Fulfillment & Source Coverage: Does the extracted_data capture the key information requested in the intent that is present in the Source Content? Are there critical omissions?
  2. 2.
    Data Accuracy: Are the values (x, y, category) and labels/title in extracted_data accurately reflecting the Source Content?
Response Format:
@@
{mdframed}
[backgroundcolor=gray!10, roundcorner=5pt] Task: Apply the suggested minor corrections to the extracted chart data.
Input:
  • Original Intent: {intent}
  • Source Content: {content}
  • Extracted Data (Pre-Refinement): {extracted_data}
  • Suggested Corrections: {suggested_corrections}
  • Expected Schema: {output_format}
Instructions:
  1. 1.
    Iterate through the Suggested Corrections.
  2. 2.
    Apply each correction to the corresponding field_path in the Extracted Data. Use suggested_value if provided, otherwise interpret the suggestion.
  3. 3.
    Ensure the final refined_data strictly follows the Expected Schema provided in the input.
  4. 4.
    Do not add new data or make changes beyond the Suggested Corrections.
Response Format:
@@
{mdframed}
[backgroundcolor=gray!10, roundcorner=5pt] Task:
@@
Input:
  • Intent: {intent}
  • Final Chart Data: {data}
Heuristics Framework:
@@
  • Time-based: $\leq$3 points $\rightarrow$ Bar; 4+ points $\rightarrow$ Line; Irregular spacing $\rightarrow$ Grouped Bar
  • Comparison: Few categories (2–5) $\rightarrow$ Bar; Many (6+) $\rightarrow$ Stacked Bar; Proportions $\rightarrow$ Pie ($\leq$6 segments)
  • Intent: Magnitude $\rightarrow$ Bar; Trend $\rightarrow$ Line; Composition $\rightarrow$ Pie/Stacked Bar
  • Anti-Patterns: Avoid cluttered pies and sparse lines. Prioritize readability.
Requirements:
  1. 1.
    Analyze the structure and nature of the Final Chart Data (types of x/y values, number of points/categories).
  2. 2.
    Relate the data structure to the Intent.
  3. 3.
    Evaluate potential chart types based on the heuristics.
  4. 4.
    Recommend the most appropriate chart type with justification.
Response Format:

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated strong capabilities in transforming text descriptions or tables to data visualizations via instruction-tuning methods. However, it is not straightforward to apply these methods directly for a more real-world use case of visualizing data from long documents based on user-given intents, as opposed to the user pre-selecting the relevant content manually. We introduce the task of intent-based chart generation from documents: given a user-specified intent and document(s), the goal is to generate a chart adhering to the intent and grounded on the document(s) in a zero-shot setting. We propose an unsupervised, two-staged framework in which an LLM first extracts relevant information from the document(s) by decomposing the intent and iteratively validates and refines this data. Next, a heuristic-guided module selects an appropriate chart type before final code generation. To assess the data accuracy of the generated charts, we propose an attribution-based metric that uses a structured textual representation of charts, instead of relying on visual decoding metrics that often fail to capture the chart data effectively. To validate our approach, we curate a dataset comprising of 1,242 <<intent, document, charts>> tuples from two domains, finance and scientific, in contrast to the existing datasets that are largely limited to parallel text descriptions/ tables and their corresponding charts. We compare our approach with baselines using single-shot chart generation using LLMs and query-based retrieval methods; our method outperforms by upto $9$ points and $17$ points in terms of chart data accuracy and chart type respectively over the best baselines.
@@ 1 Introduction
Advancements in LLMs Brown et al. (2020); Touvron et al. (2023) have enabled recent efforts in generating high-quality statistical charts from user-given text descriptions or tables Wang et al. (2023); Han et al. (2023); Maddigan and Susnjak (2023); Zhang et al. (2024); Tian et al. (2024); Zadeh et al. (2024). However, these existing approaches to automatic chart generation typically assume that users will manually provide tables or textual descriptions to be transformed into charts. Consequently, these methods cannot directly generate charts from raw documents based on specific user intents or queries. Further, many previous methods rely on instruction-tuning techniques that necessitate a substantial volume of labeled data (in the order of 10s of 1000s) for effective training. Obtaining such large corpora of charts paired with their corresponding intents is a tedious and time-consuming process, particularly for long documents (spanning >> 5 pages), which are the primary focus of our work.
@@
This paper makes four main contributions: (1) We introduce the novel task of generating charts from documents based on specific intents, to mimic real-world use cases for chart generation. (2) We propose a training-free, multi-stage framework that leverages LLMs to first execute an intent-guided iterative data extraction and refinement process from documents, and subsequently to select an appropriate chart type through a heuristic-guided deliberation, before constructing the final chart. (3) We propose a chart evaluation metric to assess the faithfulness of the generated chart data when compared to either a reference chart/ table (reference-based setting) or source document itself (reference-free setting). (4) We curate a dataset consisting of 1,242 <<intent, document, chart>> tuples from financial reports and scientific articles for this task evaluation.
@@ 2.3 Chart Evaluation
Figure 2: Sample annotations for a given SEC document and its constituent data tables. An intent and a corresponding chart are provided, along with the data table supporting the chart and a description summarizing the chart key takeaways. Other metadata such as page number and tables used from the source document are also provided. We use the page numbers to truncate the documents to include $\pm 5$ pages while using the source documents with the chart generation methods, to avoid context length issues.
@@ 3.1 Dataset Curation Overview
For the source documents, we consider two domains, namely finance and scientific. We use the U.S. Securities and Exchange Commission (SEC) 10-K filings [^1] https://www.sec.gov/edgar.shtml that are publicly available on the EDGAR website, and academic papers from *ACL conferences from the SciDuet dataset Sun et al. (2021) respectively. We scrape the HTML documents for 1,000 SEC 10-K filings, and consider all the 1,088 papers from SciDuet in the PDF form. We parse these HTML and PDF documents [^2] BeautifulSoup, https://developer.adobe.com/document-services/docs/overview/pdf-extract-api/ to obtain the tables in them separately in spreadsheets. From among the SEC filings, we first filter documents that contain table(s) with size >>7x7, and then pick the top 100 ones that have the maximum number of tables in them. As most of the numeric content that can be visualized is present in tables in these documents, we use those that densely contain them. Since academic papers do not always contain very large tables, we take the top 120 documents that have the maximum number of tables in them (and relax the size constraints). Since these documents contain multiple tables, and each table can convey several insights, we obtain multiple intents from each of them, to reduce the number of documents that are to be ingested and the overall cognitive load. Table 1 provides the source document details for the two domains. Figure 2 provides an overview of the data curation process.
@@ 5.2 Evaluation Metrics
We take inspiration in attribution as a strategy to validate specific spans of text by grounding them in the source context, both in textual space and more recently for VLMs Jiang et al. (2025); Phukan et al. (2025) We propose ChartEval, an attribution-based metric that uses structured textual representations for charts, and avoids reliance on visual decoding altogether. We trace the intermediate representations of the generated charts, extracted as structured JSONs before code generation, back to their source tables in the document(s). Each chart is represented as a collection of tuples $\langle x\text{-axis},y\text{-axis},\text{value}\rangle$ in the JSON, each of which is individually validated against the values in the corresponding tables in ground truth. For this validation, we use a modified version of the attribution algorithm from Phukan et al. (2024); Cohen-Wang et al. (2025): (i) We construct a prompt using the reference table [^4] This can be used for reference chart as well by obtaining its JSON representation, if table references are not available. as the "document", and generated chart JSON as the "output". (ii) We forward pass this prompt through the Llama-3.1-8b-Instruct model and aggregate the cross attention scores between the "output" and "document" tokens to get a token-level heatmap of <<output tokens, document tokens>> size. (iii) For each data value token in the "output", we identify the best matching span in the "document" tokens using Kadane’s algorithm Kadane (2023) on the obtained heatmap. While this formulation of ChartEval does a reference-based evaluation with the ground truth tables, it can be extended to a reference-free variant as well, where the "document" would be the entire source document context in the LLM forward pass.
@@ A.1 Qualitative examples
Task:
@@
Input: User Intent: {intent} Content: {content} Optional Feedback (if available): {optional_feedback_section} Output Format Schema: {output_format} Instructions: 1. Carefully read the User Intent. 2. Internal Thought Process (Mentally follow these steps): Decompose: Break down the intent into specific data points, labels, categories, and title. Locate: Scan the content for exact data matching the above. Extract & Structure: Collect and format data strictly according to the schema. 3. Extract relevant data points: $(x,y,\text{category})$, axis labels, and chart title. 4. If feedback is provided: Focus on fixing issues like missing elements or ignored sections. Adjust your decomposition and extraction accordingly. 5. Output must follow the JSON schema exactly. Keep numeric formats consistent. 6. Output only the JSON object. Do not include explanations or markdown like ‘‘‘json. Example Output Format:
@@
Task: Validate the extracted chart data against the source content and user intent. Determine if re-extraction is necessary or if only minor refinements are needed. Input: Original Intent: {intent} Source Content: {content} Extracted Chart Data: {extracted_data} // JSON object from the extraction step Expected Schema: {output_format} Validation Checks to Perform: 1. Intent Fulfillment & Source Coverage: Does the extracted_data capture the key information requested in the intent that is present in the Source Content? Are there critical omissions? 2. Data Accuracy: Are the values (x, y, category) and labels/title in extracted_data accurately reflecting the Source Content? Response Format:
@@
Task: Apply the suggested minor corrections to the extracted chart data. Input: Original Intent: {intent} Source Content: {content} Extracted Data (Pre-Refinement): {extracted_data} Suggested Corrections: {suggested_corrections} Expected Schema: {output_format} Instructions: 1. Iterate through the Suggested Corrections. 2. Apply each correction to the corresponding field_path in the Extracted Data. Use suggested_value if provided, otherwise interpret the suggestion. 3. Ensure the final refined_data strictly follows the Expected Schema provided in the input. 4. Do not add new data or make changes beyond the Suggested Corrections. Response Format:
@@
Task:
@@
Input: Intent: {intent} Final Chart Data: {data} Heuristics Framework:
@@
Time-based: $\leq$3 points $\rightarrow$ Bar; 4+ points $\rightarrow$ Line; Irregular spacing $\rightarrow$ Grouped Bar Comparison: Few categories (2–5) $\rightarrow$ Bar; Many (6+) $\rightarrow$ Stacked Bar; Proportions $\rightarrow$ Pie ($\leq$6 segments) Intent: Magnitude $\rightarrow$ Bar; Trend $\rightarrow$ Line; Composition $\rightarrow$ Pie/Stacked Bar Anti-Patterns: Avoid cluttered pies and sparse lines. Prioritize readability. Requirements: 1. Analyze the structure and nature of the Final Chart Data (types of x/y values, number of points/categories). 2. Relate the data structure to the Intent. 3. Evaluate potential chart types based on the heuristics. 4. Recommend the most appropriate chart type with justification. Response Format: